# MiDashengLM: Efficient Audio Understanding with General Audio Captions Horizon Team, MiLM Plus Xiaomi Inc., China

### v1

@@ 1.1 Motivation
Utilizing sound and music captions as a training target has been previously explored [1; 2; 3; 4] to enhance audio understanding.
@@
Several prominent works have demonstrated the effectiveness of ASR-based LALM training, such as Whisper [5], SpeechT5 [6], Universal Speech Model (USM) [7], Open Whisper-style Model (OWSM) [8] and Kimi-Audio [9].
@@
The training losses for ASR-based LALMs are typically low, even across different languages, suggesting that the models learn relatively little meaningful information from ASR-based data, compared to text-based training [10] (see Figure˜1).
@@ 1.2 Audio caption and speech summarization
Audio captions have been the focus of extensive research [11; 12; 13].
Most datasets during the start of the audio-caption era were manually labeled [14; 11; 12; 15], but recent work has leveraged large language models (LLMs) to scale and streamline dataset creation.
Notable LLM-assisted audio captioning datasets include WavCaps [16], AutoACD [17], SoundVECaps [18], AudiosetCaps [19] and FusionAudio-1.2M [20].
These works utilized LLMs in order to enhance existing audio captions by additional visual information [18], temporal information [17] or with additional CLAP filtering [19; 3].
@@
Neglect of spoken language: Publicly available captioning data primarily focuses on sound/music events and their audio-visual/temporal relationships, despite speech constituting the majority of real-world audio [21].
@@
Limited data diversity: Popular datasets (AudioCaps, WavCaps, AutoACD, SoundVECaps, AudiosetCaps and FusionAudio-1.2M) predominantly derive from the same audio sources (Audioset [21], VGGSound [22] and FSD50k [23]).
@@
While audio captions have been used for LALM pretraining, existing approaches typically generate new captions through either (1) paraphrasing existing descriptions [3; 24] or (2) augmenting them with (unrelated) video context [4] using LLMs, rather than genuinely diversifying the underlying audio content.
@@
General audio captions can be understood as a fusion of speech summarization [25], music captions and audio captions into one.
@@ 2 Framework
One of the most prevalent pre-trained model is the Whisper encoder [5], as seen in models like LTU-AS [26], Qwen-Audio [27], Qwen2-Audio [28], and Kimi-Audio [9], Mini-Omni [29], Llama-Omni [30], R1-AQA [31] and SALMONN [32].
Other audio encoders such as HuBERT [33], HTS-AT [34], AST [35] and BEATs [36] have also been utilized, often as secondary encoders to accommodate sound/music task knowledge.
@@
However, models such as Whisper natively does not support variable sequence lengths during training or inference and pads by default all inputs to a fixed duration of 30 seconds [28].
Changing this behavior can lead to significant performance degradation [37; 38].
@@ 3.1 ACAVCaps and Multi-Expert Chain for Audio Tasks (MECAT)
As discussed in Section˜1.1, previous captioning datasets are insufficient mainly due to the lack of speech understanding and their monotonous data source mainly stemming from Audioset [21], VGGSound [22] and FSD50k [23].
@@
We identify ACAV100M [39] as a plausible source dataset candidate for these purposes, since it has not been labeled for audio captioning before and contains little overlap with previously mentioned datasets.
@@
We began by using CED-Base [40] to predict AudioSet labels on a 2-second scale.
@@
Speech Analysis: This curation task identifies spoken language, distinguishes individual speakers, segments audio by speaker (diarization), detects speech emotion, classifies speaker gender and age and infers a transcript using Whisper [5].
@@
Having obtained all these labels, we prompt a reasoning LLM (DeepSeek-R1 [41]) in order to generate a short audio caption.
@@
MECAT is extracted from the curated dataset by filtering each source video by license and finally performing GLAP [42] to score the audio-text consistency.
@@
MECAT will also be made publicly available [43].
@@ 4 Experimental Setup
MiDashengLM is a standard Transformer-based encoder-decoder model [48], comprising a Transformer audio encoder and a text decoder.
The audio encoder builds upon Dasheng-0.6B [49], a frame-level Vision Transformer (ViT) [50] pretrained using the Masked Autoencoder (MAE) objective [51], primarily on the ACAV100M dataset.
@@
During model development, we empirically evaluated two alternative approaches: (1) integration with a frozen large language model (LLM) and (2) low-rank adaptation (LoRA [52]).
@@
The text decoder is initialized using Qwen2.5-Omni-3B [53], a publicly available pretrained language model.
@@ 5.1 Audio encoder performance
We employ the X-Ares benchmark [54], which evaluates frozen encoder embeddings through a lightweight MLP layer across three core audio domains: speech, music, and (environmental) sound.
@@ 5.2 Traditional dataset Benchmarks
We evaluate our proposed MiDashengLM on common benchmarks against two strong baselines: Qwen2.5-Omni [53] and Kimi-Audio-Instruct [9].
Note that we exclusively compare with general audio understanding models that are capable of captioning as well as spoken language understanding in order to compare fairly, since there exist work solely optimized for captions only [4; 2].
@@ 5.2.1 Audio captioning results
Results for audio captioning can be seen in Table˜7, where we select FENSE [55] as our primary audio caption metric.
@@ 5.3 Question answering results
On closed QA benchmarks (MMAU [56] and MuChoMusic [57]), MiDashengLM achieves superior performance with accuracies of 71.35% and 66.30%, respectively, outperforming all baseline models.
@@
Lastly, we evaluate MiDashengLM on our proposed MECAT-QA dataset, a part of the publicly available MECAT benchmark [43].

### v2

# MiDashengLM: Efficient Audio Understanding with General Audio Captions MiLM Plus Xiaomi Inc., China

@@ 1.1 Motivation
Utilizing sound and music captions as a training target has been previously explored [deshmukh2023pengi; kong2024audio; ghosh2024gama; ghosh2025audio] to enhance audio understanding.
@@
Several prominent works have demonstrated the effectiveness of ASR-based LALM training, such as Whisper [radford2023robust], SpeechT5 [ao2021speecht5], Universal Speech Model (USM) [zhang2023google], Open Whisper-style Model (OWSM) [peng2025owsm] and Kimi-Audio [kimi_audio].
@@
The training losses for ASR-based LALMs are typically low, even across different languages, suggesting that the models learn relatively little meaningful information from ASR-based data, compared to text-based training [touvron2023llama] (see Figure˜1).
@@ 1.2 Audio caption and speech summarization
Audio captions have been the focus of extensive research [wu2019audio; drossos2020clotho; drossos2017automated].
Most datasets during the start of the audio-caption era were manually labeled [audiocaps; wu2019audio; drossos2020clotho; martin2021ground_macs2], but recent work has leveraged large language models (LLMs) to scale and streamline dataset creation.
Notable LLM-assisted audio captioning datasets include WavCaps [mei2023wavcaps], AutoACD [sun2024autoacd], SoundVECaps [yuan2024soundvecaps], AudiosetCaps [bai2024audiosetcaps] and FusionAudio-1.2M [chen2025fusionaudio].
These works utilized LLMs in order to enhance existing audio captions by additional visual information [yuan2024soundvecaps], temporal information [sun2024autoacd] or with additional CLAP filtering [bai2024audiosetcaps; ghosh2024gama].
@@
Neglect of spoken language: Publicly available captioning data primarily focuses on sound/music events and their audio-visual/temporal relationships, despite speech constituting the majority of real-world audio [gemmeke2017audio].
@@
Limited data diversity: Popular datasets (AudioCaps, WavCaps, AutoACD, SoundVECaps, AudiosetCaps and FusionAudio-1.2M) predominantly derive from the same audio sources (Audioset [gemmeke2017audio], VGGSound [chen2020vggsound] and FSD50k [fonseca2021fsd50k]).
@@
While audio captions have been used for LALM pretraining, existing approaches typically generate new captions through either (1) paraphrasing existing descriptions [ghosh2024gama; audioflamingo3] or (2) augmenting them with (unrelated) video context [ghosh2025audio] using LLMs, rather than genuinely diversifying the underlying audio content.
@@
General audio captions can be understood as a fusion of speech summarization [retkowski2025speech], music captions and audio captions into one.
@@ 2 Framework
One of the most prevalent pre-trained model is the Whisper encoder [radford2023robust], as seen in models like LTU-AS [gong_ltuas], Qwen-Audio [qwen_audio], Qwen2-Audio [chu2024qwen2], and Kimi-Audio [kimi_audio], Mini-Omni [xie2024mini], Llama-Omni [fang2024llamaomni], R1-AQA [li2025reinforcement] and SALMONN [tang2023salmonn].
Other audio encoders such as HuBERT [hsu2021hubert], HTS-AT [chen2022htsat], AST [ast] and BEATs [chen2022beats] have also been utilized, often as secondary encoders to accommodate sound/music task knowledge.
@@
However, models such as Whisper natively does not support variable sequence lengths during training or inference and pads by default all inputs to a fixed duration of 30 seconds [chu2024qwen2].
Changing this behavior can lead to significant performance degradation [jeffries2024moonshine; liu2025voxtral].
@@ 3.1 ACAVCaps and Multi-Expert Chain for Audio Tasks (MECAT)
As discussed in Section˜1.1, previous captioning datasets are insufficient mainly due to the lack of speech understanding and their monotonous data source mainly stemming from Audioset [gemmeke2017audio], VGGSound [chen2020vggsound] and FSD50k [fonseca2021fsd50k].
@@
We identify ACAV100M [lee2021acav100m] as a plausible source dataset candidate for these purposes, since it has not been labeled for audio captioning before and contains little overlap with previously mentioned datasets.
@@
We began by using CED-Base [dinkel2023ced] to predict AudioSet labels on a 2-second scale.
@@
Speech Analysis: This curation task identifies spoken language, distinguishes individual speakers, segments audio by speaker (diarization), detects speech emotion, classifies speaker gender and age and infers a transcript using Whisper [radford2023robust].
@@
Having obtained all these labels, we prompt a reasoning LLM (DeepSeek-R1 [guo2025deepseek]) in order to generate a short audio caption.
@@
MECAT is extracted from the curated dataset by filtering each source video by license and finally performing GLAP [dinkel2025glap] to score the audio-text consistency.
@@
MECAT will also be made publicly available [mecat].
@@
Dataset
Labeling
#Vocab
Avg. Sent
Source
ClothoV2 [drossos2020clotho]
Manual
4366
11.32
Freesound
AudioCaps [audiocaps]
4844
8.70
Audioset
MusicCaps [agostinelli2023music_caps]
3730
47.17
Audioset
Songdescriber [manco2023songdescriber]
1811
26.31
MTG-Jamendo
LPMusicCaps-MTT [doh2023lpmusiccaps]
LLM
4045
25.04
MagnaTagATune
LPMusicCaps-MSD [doh2023lpmusiccaps]
14049
37.06
MillionSoundDatabase
SoundVECaps [yuan2024soundvecaps]
58401
31.48
Audioset
AutoACD [sun2024autoacd]
20491
18.47
Audioset
AudiosetCaps [bai2024audiosetcaps]
21783
28.13
Audioset + VGGSound
WavCaps [mei2023wavcaps]
24592
7.84
Audioset + BBC +
FreeSound + SoundBible
LAION-Audio-300M [laion300m]
451927
37.55
Ours‡
Reasoning-LLM
644407
22.18
ACAV100M
3.2 Training datasets and tasks
@@ 4 Experimental Setup
MiDashengLM is a standard Transformer-based encoder-decoder model [attention_is_all_you_need], comprising a Transformer audio encoder and a text decoder.
The audio encoder builds upon Dasheng-0.6B [dinkel24b_interspeech_dasheng], a frame-level Vision Transformer (ViT) [Dosovitskiy_ViT] pretrained using the Masked Autoencoder (MAE) objective [he2022masked], primarily on the ACAV100M dataset.
@@
Whisper-Large v3
Ours
Parameters
637.7M
630.3M
Pretraining data size
5M
270k
Training Objective
ASR
General captions
Context
30s
10s
Known pretraining data?
✓ [lee2021acav100m]
Open train code?
✓ [dinkel24b_interspeech_dasheng]
Open weight?
Audio-text alignment
@@
During model development, we empirically evaluated two alternative approaches: (1) integration with a frozen large language model (LLM) and (2) low-rank adaptation (LoRA [hu2022lora]).
@@
The text decoder is initialized using Qwen2.5-Omni-7B [qwen25_omni], a publicly available pretrained language model.
@@ 5.1 Audio encoder performance
We employ the X-Ares benchmark [xares_zhang], which evaluates frozen encoder embeddings through a lightweight MLP layer across three core audio domains: speech, music, and (environmental) sound.
@@ 5.2 Traditional dataset Benchmarks
We evaluate our proposed MiDashengLM on common benchmarks against two strong baselines: Qwen2.5-Omni [qwen25_omni] and Kimi-Audio-Instruct [kimi_audio].
Note that we exclusively compare with general audio understanding models that are capable of captioning as well as spoken language understanding in order to compare fairly, since there exist work solely optimized for captions only [ghosh2025audio; kong2024audio].
@@ 5.2.1 Audio captioning results
Results for audio captioning can be seen in Table˜7, where we select FENSE [fense] as our primary audio caption metric.
@@ 5.3 Question answering results
On closed QA benchmarks (MMAU [sakshi2024mmaumassivemultitaskaudio] and MuChoMusic [weck2024muchomusic]), MiDashengLM achieves superior performance with accuracies of 71.35% and 66.30%, respectively, outperforming all baseline models.
@@
Dataset
Subset
Metric
MiDashengLM
7B
Qwen2.5-Omni
7B
Kimi-Audio-Instruct
7B
MuChoMusic [weck2024muchomusic]
ACC
71.35
64.79
67.40
MMAU [sakshi2024mmaumassivemultitaskaudio]
Sound
ACC
68.47
67.87
74.17
Music
66.77
69.16
61.08
Speech
63.66
59.76
57.66
Average
66.30
65.60
64.30
MusicQA [musicqa]
FENSE
62.35
60.60
40.00
AudioCaps-QA [wang2024audiobench]
54.31
53.28
47.34
5.3.1 MECAT-QA
Lastly, we evaluate MiDashengLM on our proposed MECAT-QA dataset, a part of the publicly available MECAT benchmark [mecat].
@@ 6 Conclusion
Appendix A Data sources
A.1 Speech datasets
Table 14: Speech training data. The notation † leverages Whisper to generate automatic transcripts by the authors. The column “SFT ?” indicates whether the dataset is used for supervised finetuning. By default all data is used for pretraining.
Table 15: General Sound and Audio Datasets. ACAVCaps is utilized for audio-text alignment. The column “SFT ?” indicates whether the dataset is used for supervised finetuning. By default all data is used for pretraining.
Dataset
Task
Length (h)
SFT ?
FSD50k [fonseca2017freesound]
Sound Event
77
AudioSet [gemmeke2017audio]
5,200
AudioSet-strong [audioset_strong]
220
VGGSound [chen2020vggsound]
540
FSDKaggle2018 [fonseca2018generalfsd]
20
FSDKaggle2019 [fonseca2019audio]
100
ARCA23k [arca23k]
120
AutoACD [sun2024autoacd]
Audio (Sound) Caption
5,200
AudioSetCaps [bai2024audiosetcaps]
6,000
SoundVECaps [yuan2024soundvecaps]
5,000
WavCaps [mei2023wavcaps]
7,567
Audiocaps [audiocaps]
100
Clothov2 [drossos2020clotho]
17
TACOS [primus2025tacos]
98
CochlScene [jeong2022cochlscene]
SoundScape
500
BirdSet [birdset]
7,000
ACAVCaps
General Caption
38,662
Total
76,421
69,081
A.3 Speech and paralinguistic datasets
Table 16: Speech and sound paralinguistic datasets. The column “SFT ?” indicates whether the dataset is used for supervised finetuning. By default all data is used for pretraining.
Dataset
Task
Length (hours)
SFT ?
IEMOCAP [busso2008iemocap]
Emotion
8
Meld [poria2018meld]
12
SUBESCO [sultana2021subesco]
9
RAVDESS-Speech [livingstone2018ravdess]
2
RAVDESS-Song [livingstone2018ravdess]
1
CREMA-D [cao2014crema]
4
ESD [zhou2021esd]
29
VocalSound [gong2022vocalsound]
Vocal Sound classification
20
NonSpeech7k [rashid2023nonspeech7k]
3
VoxLingua107 [valk2021voxlingua107]
Language Identification
7,200
CommonLanguage [commonlanguage]
45
YLACombe [ylacombe]
5
VoxCeleb1 [nagrani17_interspeech]
Speaker verification
76
CNCeleb [fan2020cnceleb]
Speaker verification
2,100
VoxCeleb2 [chung2018voxceleb2]
Speaker age
Speaker verification
Gender classification
1,000
VoxBlink1 [lin2024voxblink]
Speaker verification
1,300
VoxBlink2 [lin2024voxblink2]
Speaker verification
2,600
VoxTube [yakovlev2023voxtube]
Speaker verification
Language Identification
Gender classification
5,200
LibriCount [libricount]
Speaker counting
8
FluentSpeechCommands [lugosch2019speech]
Intent Classification
17
speechocean762 [zhang2021speechocean762]
Gender
Speaker age
5
ASVSpoof5 [wang2024asvspoof]
Spoof detection
603
Total
20,247
19,572
A.4 Music Datasets
Table 17: Music-Related Datasets Overview. The column “SFT ?” indicates whether the dataset is used for supervised finetuning. By default all data is used for pretraining.
Dataset
Task
Length (h)
SFT ?
MusicCaps [agostinelli2023music_caps]
Music Caption
15
Songdescriber [manco2023songdescriber]
23
LPMusicCaps-MTT [doh2023lpmusiccaps]
18
LPMusicCaps-MSD [doh2023lpmusiccaps]
1,000
VocalSet [wilkins2018vocalset]
Singing style identification
10
FreeMusicArchive [nsynth2017]
Genre recognition
610
MTG-Jamendo [bogdanov2019mtgjamendo]
Instrument classification
Genre recognition
3,768
NSynth [nsynth2017]
Instrument classification
360
GoodSounds [bandiera2016good]
28
chMusic [gong2021chmusic]
1
CTIS [CTIS]
1
Total
5,824
5,814
A.5 Question Answering Datasets
Table 18: Question answering datasets used in this work. Datasets denoted with † have been modified from their original dataset by using an LLM to change captions into question-answer pairs. We display the number of questions and answers in each dataset as # QA. The column “SFT ?” indicates whether the dataset is used for supervised finetuning. By default only AVQA, MusicQA ad ClothoAQA are used during pretraining.